OpenStudy (anonymous):

A rectangular garden has a perimeter of 92 feet. If the length of the garden is 1 foot greater than twice its width, what is the length of the garden in feet?

OpenStudy (lilg132):

length = 2x + 1 width = x perimeter = 2(2x + 1) + 2x 92 = 2(2x + 1) + 2x 92 = 4x + 2 + 2x 92 - 2 = 6x 90 = 6x 90 / 6 = 6x / 6 15 = x width = 15 length = 2(15 + 1) length = 30 + 1 length = 31 perimeter = 2(31) + 2(15) perimeter = 62 + 30 perimeter = 92 remember to click good answer

OpenStudy (anonymous):

great, thanks a lot.